OpenStudy (anonymous):

Compute the Gross Pay and then the Net Pay for the following situations 4. Total Hours Worked: 42 Regular Pay Rate: $8.00/hour Federal Tax Withheld: $38.00 FICA Withheld: 6.2% Medicare Withheld: 1.45% Georgia Tax Withheld: $12.00 Health Insurance: $15.00 401K Contribution: $20.00

OpenStudy (anonymous):

@sammixboo

OpenStudy (anonymous):

do you know how to do this ma'am ? @sammixboo

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

@Mehek14

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

Gross Pay = (# of hours worked)*(wage) wage = amount paid per hour

jimthompson5910 (jim_thompson5910):

Gross pay is the amount of money BEFORE taxes and other things are deducted from the pay Net pay is the amount of money AFTER taxes and other deductions

OpenStudy (anonymous):

can you help me with the first part please like start me off @jim_thompson5910

jimthompson5910 (jim_thompson5910):

what did you get for the gross pay

OpenStudy (anonymous):

how do i get the gross pay it looks confusing like what is the formul ? @jim_thompson5910

jimthompson5910 (jim_thompson5910):

I gave it to you Gross Pay = (# of hours worked)*(wage)

jimthompson5910 (jim_thompson5910):

take the number of hours worked and multiply it by the wage

jimthompson5910 (jim_thompson5910):

of course, this isn't considering overtime but we won't worry about that

OpenStudy (anonymous):

so i multiply 42 by 8 ?

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

correct

OpenStudy (anonymous):

which is 336

OpenStudy (anonymous):

sir @jim_thompson5910

jimthompson5910 (jim_thompson5910):

correct

jimthompson5910 (jim_thompson5910):

gross pay = $336

jimthompson5910 (jim_thompson5910):

All of this shown below `Federal Tax Withheld: $38.00` `FICA Withheld: 6.2%` `Medicare Withheld: 1.45%` `Georgia Tax Withheld: $12.00` `Health Insurance: $15.00` `401K Contribution: $20.00` will be the taxes and deductions taken out of the gross pay

jimthompson5910 (jim_thompson5910):

do you see how to add up all that?

OpenStudy (anonymous):

yes i have to add all of them to get the net pay sir ? @jim_thompson5910

jimthompson5910 (jim_thompson5910):

so what do you get when you add them all up (ignore the percentages for now)

OpenStudy (anonymous):

i got 85.765 is that correct sir ?

jimthompson5910 (jim_thompson5910):

ignore the percentages for now

jimthompson5910 (jim_thompson5910):

Federal Tax Withheld: $38.00 Georgia Tax Withheld: $12.00 Health Insurance: $15.00 401K Contribution: $20.00

OpenStudy (anonymous):

70 sir @jim_thompson5910

jimthompson5910 (jim_thompson5910):

no

OpenStudy (anonymous):

85

jimthompson5910 (jim_thompson5910):

Add up these Federal Tax Withheld: $38.00 Georgia Tax Withheld: $12.00 Health Insurance: $15.00 401K Contribution: $20.00

OpenStudy (anonymous):

i got 85 sir

jimthompson5910 (jim_thompson5910):

now we consider the percentages

jimthompson5910 (jim_thompson5910):

FICA Withheld: 6.2% Medicare Withheld: 1.45% 6.2% + 1.45% = 7.65%

jimthompson5910 (jim_thompson5910):

we will take 7.65% of the gross pay to get 0.0765*336 = 25.704 = 25.70 That means we add 25.70 to 85 to get 25.70+85 = 110.70

jimthompson5910 (jim_thompson5910):

$110.70 is the total amount you take out of the gross pay to get the net pay Net Pay = (Gross Pay) - (Total Deductions) Net Pay = (336) - (110.70) Net Pay = 225.30

OpenStudy (anonymous):

omg thank you so much you are amazing :)

jimthompson5910 (jim_thompson5910):

np

OpenStudy (anonymous):

would you mind helping me with others ? @jim_thompson5910

jimthompson5910 (jim_thompson5910):

sure, a few more

jimthompson5910 (jim_thompson5910):

post it as a new question though

OpenStudy (anonymous):

yes i will @jim_thompson5910